# ATPrompt: Textual Prompt Learning with Embedded Attributes

### v1

@@ 7 Additional Experiments
7.1 Domain Generalization
Tab. 8 shows the domain generalization results for three baseline methods. The results show that our method improves CoOp, CoCoOp and MaPLe methods by 0.90%, 0.49% and 0.33% respectively.
@@ 7.2 Ablation Study
Attribute Order. We have conducted some experiments in the main paper and verified that the order of attributes does not significantly impact model performance. The fluctuation of the results is within a reasonable range. Due to page limitations, we only provide a portion of the results in the main paper. In Tab. 9, we conduct more experiments to verify our observations.
@@
Attribute Position. In addition to assessing the impact of the class token’s position within the textual prompt on performance, we also examine the influence of the attribute token’s positioning. Fig. 10 and Tab. 10 present the visualization of attribute positions and the corresponding experimental results, respectively. The results indicate that the interval version achieves the best performance among all variations.
@@
Initialization. Existing baseline methods initialize the soft tokens using the embeddings of the phrase “a photo of a” in their official implementations across 11 recognition datasets. However, in our approach, the addition of attribute tokens may render this initialization strategy inappropriate.
In our method, the class soft tokens $[T_{1}],...,[T_{M}]$ are randomly initialized by sampling from a zero-mean Gaussian distribution with a standard deviation of 0.02. In Tab. 11, we compare the effects of different initialization methods on model performance. Experimental results show that random initialization provides a superior starting point for training.
Attribute Base Novel HM
“a photo of a” 76.40 70.07 73.10
Random Normal Init 76.27 70.60 73.33
8 Discussion
8.1 Comparison with Querying LLM Directly.
There are two disadvantages to directly asking a large language model to obtain universal attributes. First, determining the optimal content of these attributes is challenging. Second, identifying the ideal number of attributes is equally difficult. However, if a user wishes to acquire attributes directly in this manner, it remains a feasible approach. Our experiments have shown that the optimal number of attributes is often around two. Therefore, if the user prefers to avoid the attribute search process, directly asking the language model to summarize two universal attributes for known categories and then using them for training is a more convenient method, albeit at the cost of some performance.
9 Limitations and Future Works.
(1) Our proposed approach utilizes a differentiable search method that learns to identify suitable attributes for our prompt format. While this method avoids lengthy training times and high costs, there is still room for optimization. We attempted to employ a Large Language Model (LLM) to discover suitable attributes, but determining the optimal number of attributes for the current format remains a challenge. In future research, we plan to investigate how to leverage the Chain-of-Thought (CoT) method to further explore the potential of Multimodal Large Language Models (MLLM) in enhancing the attribute discovery process. (2) Our approach involves embedding fixed explicit attributes into soft prompts to enhance model performance. In the future, we intend to explore transitioning from using explicit fixed attributes to implicit, learnable attributes. This shift aims to enable the model to autonomously discover suitable attributes in a learning-driven manner, thereby improving model performance.
@@ 8 Discussion
Combinations & Weights (shape), score: 0.298 (color), score: 0.004 (material), score: 0.002 (function), score: 0.002 (size), score: 0.003 (shape, color), score: 0.003 (shape, material), score: 0.006 (shape, function), score: 0.000 (shape, size), score: 0.565 (color, material), score: 0.000 (color, function), score: 0.001 (color, size), score: 0.005 (material, function), score: 0.000 (material, size), score: 0.002 (function, size), score: 0.002 (shape, color, material), score: 0.002 (shape, color, function), score: 0.002 (shape, color, size), score: 0.000 (shape, material, function), score: 0.001 (shape, material, size), score: 0.085 (shape, function, size), score: 0.001 (color, material, function), score: 0.001 (color, material, size), score: 0.000 (color, function, size), score: 0.002 (material, function, size), score: 0.001 (shape, color, material, function), score: 0.001 (shape, color, material, size), score: 0.001 (shape, color, function, size), score: 0.001 (shape, material, function, size), score: 0.005 (color, material, function, size), score: 0.001 (shape, color, material, function, size), score: 0.001
Table 13: Output results after 40 epochs of training on Caltech101.

### v2

@@ 7.1 Ablation Study
Attribute Order. In the main paper, our experiments confirm that the order of attributes does not significantly impact model performance, with results fluctuating within an acceptable range. Here we provide additional experiments in Tab. 9 to support this observation.
@@
(shape, color) 76.32 70.39 73.24
(color, shape) 76.27 70.60 73.33
(size, habitat) 76.44 70.23 73.20
(habitat, size) 76.46 70.16 73.14
@@
Attribute Position. We also investigated the impact of attribute token positioning within the prompt. Fig. 10 visualizes the positions tested, and Tab. 10 presents the results. Our findings show that the ”interval” configuration, where attributes are placed between class tokens, yields the best performance.
@@
Initialization. Baseline methods typically initialize soft tokens using the embeddings of the phrase ”a photo of a.” The inclusion of attribute tokens makes this strategy suboptimal for our method. We instead initialize class soft tokens ($[T_{1}],...,[T_{M}]$) by sampling from a Gaussian distribution with a mean of 0 and a standard deviation of 0.02. As shown in Table 12, this random initialization provides a superior starting point for training.
Attribute Base Novel HM
“a photo of a” 76.40 70.07 73.10
Random Normal Init 76.27 70.60 73.33
8 Discussion
Comparison with Direct LLM Queries. Directly querying an LLM for universal attributes presents two challenges: determining the optimal attribute content and identifying the ideal number of attributes. Our experiments suggest that two attributes are often optimal. Therefore, users can bypass our search process by directly prompting the LLM to summarize two universal attributes. This offers a simpler approach, though it may result in a slight performance trade-off.
Why do attributes searched on a source dataset (ImageNet) generalize well? The attributes identified on ImageNet (e.g., color, shape) are fundamental properties of natural objects. Representations learned under the guidance of these universal attributes are therefore inherently generalizable and transfer effectively to other datasets and classes.
Why does ATPrompt not outperform regularization-based methods in isolation? ATPrompt is a plug-in module designed to optimize the prompt’s structure. In contrast, regularization-based methods are often comprehensive frameworks that employ multiple components (e.g., learnable visual prompts, MLPs) simultaneously. While ATPrompt may not outperform these multi-faceted approaches on its own, its strength lies in its ability to be integrated into other methods, consistently improving their performance beyond previous baselines.
9 Limitations and Future Works.
Beyond the limitations discussed in the main paper, we identify the following directions for future research: (1) While our differentiable search method is efficient, we aim to further enhance the attribute discovery process. A promising direction is to leverage Multimodal Large Language Models (MLLMs), potentially using techniques like Chain-of-Thought (CoT), to better automate the selection of optimal attribute content and quantity. (2) Our current approach embeds fixed, explicit attributes into the prompt. In the future, we plan to explore a transition to implicit, learnable attributes. This would enable the model to discover optimal attributes in a data-driven manner during training, potentially unlocking further performance gains.
@@ 8 Discussion
Attribute Combinations & Corresponding Weights (shape), weight: 0.298 (color), weight: 0.004 (material), weight: 0.002 (function), weight: 0.002 (size), weight: 0.003 (shape, color), weight: 0.003 (shape, material), weight: 0.006 (shape, function), weight: 0.000 (shape, size), weight: 0.565 (color, material), weight: 0.000 (color, function), weight: 0.001 (color, size), weight: 0.005 (material, function), weight: 0.000 (material, size), weight: 0.002 (function, size), weight: 0.002 (shape, color, material), weight: 0.002 (shape, color, function), weight: 0.002 (shape, color, size), weight: 0.000 (shape, material, function), weight: 0.001 (shape, material, size), weight: 0.085 (shape, function, size), weight: 0.001 (color, material, function), weight: 0.001 (color, material, size), weight: 0.000 (color, function, size), weight: 0.002 (material, function, size), weight: 0.001 (shape, color, material, function), weight: 0.001 (shape, color, material, size), weight: 0.001 (shape, color, function, size), weight: 0.001 (shape, material, function, size), weight: 0.005 (color, material, function, size), weight: 0.001 (shape, color, material, function, size), weight: 0.001
Table 13: Output results after 20 epochs of attribute searching on the Caltech101 dataset.